OpenStudy (anonymous):

-28xy over 4xSquared

OpenStudy (anonymous):

\[ \frac{-28xy}{4^2} = \frac{-7xy}{4} \]

OpenStudy (anonymous):

-28xy over 4x^2

OpenStudy (anonymous):

\[ -\frac{7y}{x} \]

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

so is it -7y over x?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Absolutely positive